### v2

@@ 1.2 Applications
A Scaling technique with Applications to Random Walks and PageRank.
Let $\mathbf{A}\in\mathbb{R}^{n\times n}$ be the adjacency matrix of an $n$-node graph and let $\mathbf{D}\in\mathbb{R}^{n\times n}$ be the diagonal matrix with vertex-degrees on the diagonal. Then the normalized Laplacian is $\widetilde{\mathbf{L}}=\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$. Even though $\widetilde{\mathbf{L}}$ is not Boolean, we can still use Theorem 1.1 since $\widetilde{\mathbf{L}}v=\mathbf{D}^{-1/2}(w-\mathbf{A}w)$ for $w=\mathbf{D}^{-1/2}v$ and the product $\mathbf{A}w$ can be computed via Theorem 1.1. The same technique extends to random-walk probability matrices. In turn, this allows fast simulations of Markov chains to compute stationary distributions and iterative computation of the PageRank (Page et al., 1999) via Theorem 1.1.
Complexity-Theoretic Implications.
The Boolean matrix multiplication (BMM) conjecture states that no algorithm can multiply two boolean $n\times n$ matrices in $O(n^{3-\epsilon})$ time for constant $\epsilon>0$ (Basch et al., 1995; Williams and Williams, 2010). This conjecture is used to prove lower bounds on combinatorial algorithms (i.e., algorithms that do not make use of Strassen-style fast matrix multiplication methods), motivated by the property that fast matrix multiplication, while being theoretically fast (currently $O(n^{2.372})$ Williams et al. (2024)), is very slow in practice due to large constants hidden in $O$-notation. Hence BMM gives a lower bound on more practical algorithms and can be used to prove that any further complexity improvement would require fast matrix multiplication.
Theorem 1.1 implies that for matrices with corrupted VC-dimension $d$, BMM can be solved in $\widetilde{O}(n^{3-1/d})$ time, allowing for improved combinatorial upper bounds. The BMM problem was the original motivation for Gasieniec and Lingas (2003). Moreover, as demonstrated in Alves et al. (2024), this algorithm is also highly efficient in practice and does not encounter the large constant of fast matrix multiplication. For matrices of VC-dimension $1$, Theorem 1.1 achieves the optimal matrix-multiplication time of $\widetilde{O}(n^{2})$ (upto polylogarithmic factors). Through this view, Theorem 1.1 implies that problems with lower bounds based on BMM, may be solved more efficiently on structured inputs.
@@
Suppose we are given Boolean matrices $\mathbf{M}_{1},\mathbf{M}_{2}\in\{0,1\}^{n\times n}$, such that $\mathbf{M}_{i}$ has corrupted VC-dimension $d_{i}$ and $\mathbf{M}_{i}^{\top}$ has corrupted VC dimension $d_{i}^{\prime}$ for $i\in\{1,2\}$. Let $s=\min\{d_{1},d_{2},d_{1}^{\prime},d_{2}^{\prime}\}$. Then, there is an algorithm for boolean matrix multiplication with in $O(n^{3-\frac{1}{s}})$ time complexity.
@@
This conjecture is used to prove conditional lower bounds for many dynamic problems, such as maintaining shortest paths, effective resistances, reachability, maximum flow, bipartite matching, triangle detection, and many others in graphs undergoing vertex insertions and deletions (Henzinger et al., 2015). Even with use of fast matrix multiplication, no algorithm can beat $O(n^{2})$ time on dense graphs, which for many of these problems is equivalent to re-solving them from scratch whenever the graph changes. Since Theorem 1.1 shows the OMv conjecture does not hold for structured matrices, it opens up improved dynamic algorithms.
We next list some examples of new dynamic algorithms obtained via this technique, whose upper bounds are only achievable due to their structured nature, otherwise violating either the BMM or OMv conjecture.
Implications for Dynamic Algorithms.
In the following, we state results for dynamic graph algorithms. These are data structures that maintain some graph property while the underlying graph changes over time. We defer the formal proofs for these results to Appendix B.
An important class of instances of linear equations that arise in practice has the form $\mathbf{L}x=b$, where $\mathbf{L}$ is the Laplacian of an undirected graph $G=(V,E)$ given by $\mathbf{L}\coloneqq\mathbf{D}-\mathbf{A}$ where $\mathbf{D}$ is a diagonal matrix such that $\mathbf{D}_{i,i}=\mathrm{deg}(i)$ for $i\in[n]$ where $|V|=n$, and $\mathbf{A}$ is the adjacency matrix of the graph. There are Laplacian solvers that run in time nearly linear in the number of edges of the underlying graph (Spielman and Teng, 2004). We study the problem of constructing a dynamic Laplacian solver for $\mathbf{L}_{t}x=b_{t}$ where the Laplacian changes over time through vertex updates to the graph, and each $b_{t}$ is an arbitrary new vector.
Previous upper bounds could only handle a $\operatorname{poly}(1/\epsilon)$ accuracy by maintaining spectral sparsifiers (Abraham et al., 2016; Bernstein et al., 2022) or vertex sparsifiers (Durfee et al., 2019; Chen et al., 2020). Moreover, exact or $O(\log 1/\epsilon)$-time dependencies are ruled out by the OMv conjecture (since no algorithm beats naive recomputation from scratch in $\widetilde{O}(|E|)=\widetilde{O}(n^{2})$ time via nearly-linear time Laplacian system solvers (Spielman and Teng, 2004)). We provide the first result faster than naive recomputation from scratch in the high-accuracy $\log(1/\epsilon)$ regime.
@@
Using the fast dynamic Laplacian solver, we maintain dynamic effective resistances of a graph, which is a critical computation for graph clustering (Alev et al., 2017), fault-tolerant computing (Ghosh et al., 2008), network analysis (Tizghadam and Leon-Garcia, 2008; Anand and Umans, 2023; Anand et al., 2024a), and biological systems (Klein et al., 2022), as it signifies the degree of connectivity between two vertices. The effective resistance $r_{G}(a,b)$ between vertex $a$ and vertex $b$ in a graph $G$ is a graph-theoretic analog of leverage scores that represents the energy needed to route one unit of electric flow from $a$ to $b$, and has a natural linear algebraic formulation: let $\mathbf{e}_{i}$ denote the $i$-th standard basis vector. Then, for all $a,b\in V$, $r_{G}(a,b)=(\mathbf{e}_{a}-\mathbf{e}_{b})^{\top}\mathbf{L}^{\dagger}(\mathbf{e}_{a}-\mathbf{e}_{b})$, where $\mathbf{L}$ is the Laplacian of $\mathbf{G}$. Finally, $\mathbf{L}^{\dagger}$ denotes the Moore-Penrose pseudoinverse of $\mathbf{L}$.
@@
(Dynamic Effective Resistance). There is a dynamic algorithm that, given a dynamic graph $G=(V,E)$ with corrupted VC-dimension bounded by $d$, maintains effective resistances in $G$. The data structure supports queries that receive a pair of vertices $u,v\in V$ and error parameter $\epsilon>0$. Then, in $\widetilde{O}(n^{2-1/d}\log 1/\epsilon)$ time, the algorithm returns a $(1\pm\epsilon)$-approximation of the effective resistance. Moreover, each node update to $G$ in the dynamic data structure takes $\widetilde{O}(n)$ time.
Our dynamic matrix-vector multiplication data structure also leads to faster dynamic triangle detection algorithms for graphs with corrupted VC-dimension $d$, which are critical components of fraud detection systems and community-detection algorithms. No subquadratic time algorithm exists for vertex updates, conditional on BMM (Abboud and Vassilevska Williams, 2014), and also conditional on OMv (Henzinger et al., 2015). We show that these lower bounds can be broken on structured graphs.
@@
(Dynamic Triangle Detection). There is a dynamic algorithm that, given a dynamic graph $G=(V,E)$ with corrupted VC-dimension $d$, maintains whether $G$ contains a triangle or not. Each vertex update to $G$ takes $O(n^{2-1/d})$ time and returns a Boolean indicator for whether $G$ contains a triangle or not.
Furthermore, our techniques lead to faster dynamic approximate algorithms for single-source distances. In this problem, we are tasked with finding the distances from a designated source vertex to every other vertex.
@@
(Dynamic Approximate Single-Source Shortest Paths). There is a dynamic algorithm that maintains $(1+\epsilon)$-approximate single source distances on a dynamic unweighted undirected graph $G=(V,E)$. If the corrupted VC-dimension of $G$ is bounded by $d$, each vertex update to $G$ takes $\widetilde{O}(kn^{2-1/(2d)}/\epsilon)$ time, and querying the distances for any source node takes $O(n^{2-1/(2d)}/\epsilon)$ time.
Given a metric space $\mathcal{M}$ with $n$ points and a positive integer $k\leq n$, the $k$-center problem asks one to select $k$ points, referred to as centers, such that the maximum distance of any point in the metric space to its closest center is minimized. It is known that the $k$-center is NP-hard to approximate within a factor of $(2-\epsilon)$ for any $\epsilon>0$ (Hsu and Nemhauser, 1979). The best existing dynamic algorithms (Cruciani et al., 2024) only work with edge updates and use fast matrix multiplication as a black-box. We present an algorithm without fast matrix multiplication that supports more powerful vertex updates.
@@
(Dynamic Approximate $k$-center). Given an unweighted undirected graph $G=(V,E)$, there is a dynamic algorithm for $(2+\epsilon)$-approximate $k$-center with node update time $O(kn^{2-1/2d}/\epsilon)$, where $d$ is a bound on the corrupted VC-dimension of $G$.
@@ 1.3 Further Related Work
Results for general Boolean matrix multiplication.
A line of work shows that the complexity of matrix-vector multiplication for general Boolean matrices (where the multiplication is over the Boolean semiring) can attain a subpolynomial speed-up. For instance, Williams (2007) showed that BMM can be computed in $O(n^{3}/\log n)$ time. Larsen and Williams (2017) extended this to $O(n^{3}/2^{\Omega(\sqrt{\log n})})$. This was further extended in Abboud et al. (2024) to $O(n^{3}/2^{\Omega(\sqrt[7]{\log n})})$.
Results for structured matrices.
When the underlying matrix is Vandermonde, Toeplitz, Hankel, or Cauchy, a body of work shows that they admit fast matrix-vector multiplications through convolutional transformations (Motzkin, 1951; Olshevsky and Shokrollahi, 2000; Pan and Tsigaridas, 2014). Orthogonal-vector (OV) matrices are structured matrices where each row and column receives a label $v\in\{0,1\}^{d}$ and $\mathbf{M}_{i,j}=1$ if and only if the corresponding two labels are orthogonal. Similar to our corrupted VC-dimension results, Alman and Vassilevska Williams (2020) show that such OV matrices allow for fast matrix-vector products if there are few corruptions to the matrix.
Another approach to handle structured inputs is through the algorithms with predictions regime, where some learning algorithm extracts the structure and can predict information about future vectors, leading to faster matrix-vector multiplication algorithms Henzinger et al. (2024); Brand et al. (2024).
Result for dynamic structured graphs.
Our dynamic algorithms hold for any structured graph, even if we do not know their structure. When the structure is known, there exist specialized dynamic algorithms tailored to the specific graph class. These include, for instance, interval graphs Crespelle (2019); Chen et al. (2024), geometric graphs defined from Kernels (Alman et al., 2020), planar graphs (Korhonen et al., 2024), dynamic planar graphs (Abboud and Dahlgaard, 2016), and H-minor free graphs (Dorn et al., 2012).
Matrix vector multiplication in optimization.
For convex optimization algorithms, there is a long history on developing data structures that facilitate fast matrix-vector products. Here the matrices are generally projection matrices that require computation of some matrix inverse. While historically, most research was on accelerating the maintenance of the matrix inverse (Karmarkar, 1984; Vaidya, 1989; Lee and Sidford, 2015; Cohen et al., 2021; Lee et al., 2019; Brand, 2020; Jiang et al., 2021, 2020a; Huang et al., 2022; Jiang et al., 2022; Lee et al., 2015; Jiang et al., 2020b), this research has progressed so far that the current bottleneck for the fastest linear program solvers are simple matrix-vector products (Brand et al., 2020, 2021).
@@ 2 Preliminaries
Let $[n]=\{1,\dots,n\}$. We use $O(\cdot)$ to hide constant factors, and $\widetilde{O}$ to hide polylogarithmic factors. Let $\|\cdot\|_{1}$ denote the $\ell_{1}$ (Manhattan distance). For any matrix $\mathbf{M}$, let $\mathbf{M}_{i}$ and $\mathbf{M}_{:,i}$ denote the $i$’th column and $i$’th row (respectively) of $\mathbf{M}$. For a set $V$ we write $\mathbf{M}\in\mathbb{R}^{m\times V}$ for the $m\times|V|$ matrix where we can index columns by $x\in V$, i.e., $\mathbf{M}_{x}\in\mathbb{R}^{m}$. For any vector $x\in\mathbb{R}^{n}$, the Hamming weight of $x$ is the number of non-zero entries of $x$ also denoted by $\operatorname{nnz}(x)$. For Boolean vectors $x,y\in\{0,1\}^{n}$, the Hamming distance between $x$ and $y$ is given by $\|x-y\|_{1}$. Finally, for any vector $x\in\mathbb{R}^{n}$ and positive definite matrix $\mathbf{L}\in\mathbb{R}^{n\times n}$, define the $\mathbf{L}$-induced norm of $x$ by $\|x\|_{\mathbf{L}}=\sqrt{x^{\top}\mathbf{L}x}$.
VC-Dimension of Range Spaces.
A range space (also called a set system) is a pair $\mathfrak{R}=(X,R)$ where $X$ is a set and $R$ is a set of subsets of $X$. Let $\Pi_{R}(A)=\{A\cap r:r\in R\}$. The VC-dimension of $\mathfrak{R}$ is given by $\mathrm{VC}(\mathfrak{R})=\max\{|S|:S\subseteq X\text{ and }|\Pi_{R}(S)|=2^{|S|}\}$. The primal shatter function $\pi$ of $\mathfrak{R}$ is given by $\pi(m)=\max\{|\Pi_{R}(A)|:A\subseteq X,|A|=m\}\text{ for $m\geq 0$}$. Let $B$ be a set of ranges in $R$. Then, let $\Pi_{X}^{*}(B)=\{C:C\subseteq X\text{and $C$ is not crossed by any range in $Q$}\}$ and let $\Pi_{X}^{*}(B)$ be maximal. Then, for $m\geq 0$, the dual shatter function $\pi^{*}$ of the range space is given by $\pi^{*}(m)=\max\{|\Pi_{X}^{*}(B)|:B\subseteq R,|B|=m\}$.
VC-Dimension of Matrices.
@@
Any matrix $\mathbf{M}\in\{0,1\}^{m\times n}$ corresponds to a set family given by $\mathcal{F}_{\mathbf{M}}=\{\{j:\mathbf{M}_{i,j}=1\}:i\in[m]\}$, i.e., consider each row of $\mathbf{M}$ as the indicator vector of some set. Through it, we can encode $\mathbf{M}$ by the range space $\mathfrak{R}_{M}=([n],\mathcal{F}_{\mathbf{M}})$. Then, the VC-dimension of $\mathbf{M}$ is defined as the VC-dimension of $\mathfrak{R}_{\mathbf{M}}$. Here, a subset $S$ of the columns of $\mathbf{M}$ is said to be shattered if each of the $2^{|S|}$ many 0/1 strings appears in some row in the restriction of $\mathbf{M}$ to the columns in $S$. Then, the VC-dimension of $\mathbf{M}$ is the maximum size of a shattered subset of the columns of $\mathbf{M}$. Moreover, the VC-dimension $d_{m,n}$ of a matrix $\mathbf{M}\in\{0,1\}^{m\times n}$ satisfies $1\leq d_{m,n}\leq\log n$.
Pollard Pseudodimension of Matrices.
For a family of functions $\mathcal{F}$ mapping $V$ to $\{0,\dots,\ell-1\}$, the Pollard pseudodimension of $\mathcal{F}$, denoted by $\mathrm{Pdim}(\mathcal{F})$, is the largest integer $d$ such that there exists $x_{1},\dots,x_{d}\in V$ and thresholds $y_{1},\dots,y_{d}\in[\ell]$ such that for any $d$-length bit vector $\vec{b}=(b_{1},\dots,b_{d})\in\{0,1\}^{d}$, there exists an $f\in\mathcal{F}$ such that for all $i\in[d],f(x_{i})\leq y_{i}$ if and only if $b_{i}=1$. For a real matrix $\mathbf{M}$, define the pseudo-dimension of $\mathbf{M}$, denoted by $\mathrm{Pdim}(\mathbf{M})$ by thinking of the rows of $\mathbf{M}$ as functions and taking the pseudo-dimension of the resulting class of functions. Specifically, if $\mathbf{M}$ is an $m\times n$ matrix, for each $i\in\{1,\dots,m\}$, define $f_{\mathbf{M},i}:\{1,\dots,n\}\to\mathbb{R}$ by $f_{\mathbf{M},i}(j)=\mathbf{M}_{i,j}$, and let $\mathrm{Pdim}(\mathbf{M})=\mathrm{Pdim}(\{f_{\mathbf{M},i}:i\in\{1,\dots,m\}\})$.

### v3

@@ 1.2 Applications
Random Walks and PageRank by Scaling. Let $\mathbf{A}\in\mathbb{R}^{n\times n}$ be the adjacency matrix of an $n$-node graph and let $\mathbf{D}\in\mathbb{R}^{n\times n}$ be its degree matrix. The normalized Laplacian is $\widetilde{\mathbf{L}}=\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$. Even though $\widetilde{\mathbf{L}}$ is not Boolean, we can still use Theorem˜1.3 since $\widetilde{\mathbf{L}}v=\mathbf{D}^{-1/2}(w-\mathbf{A}w)$ for $w=\mathbf{D}^{-1/2}v$ and the product $\mathbf{A}w$ can be computed via Theorem˜1.3. The same technique extends to random-walk probability matrices. In turn, this allows fast simulations of Markov chains to compute stationary distributions and iterative computation of the PageRank (Page et al., 1999) via Theorem˜1.3.
Complexity-Theoretic Implications. The Boolean matrix multiplication (BMM) conjecture states that no algorithm can multiply two boolean $n\times n$ matrices in $O(n^{3-\epsilon})$ time for constant $\epsilon>0$ (Basch et al., 1995; Williams and Williams, 2010). This conjecture is used to prove lower bounds on combinatorial algorithms (i.e., algorithms that do not make use of Strassen-style fast matrix multiplication methods), motivated by the property that fast matrix multiplication, while being theoretically fast (currently $O(n^{2.372})$ Williams et al. (2024)), is very slow in practice due to large constants hidden in $O$-notation. Hence BMM gives a lower bound on more practical algorithms and can be used to prove that any further complexity improvement would require fast matrix multiplication.
Theorem˜1.3 implies that for matrices with corrupted VC-dimension $d$, BMM can be solved in $\widetilde{O}(n^{3-1/d})$ time, allowing for improved combinatorial upper bounds. The BMM problem was the original motivation for Björklund and Lingas (2001); Gasieniec and Lingas (2003). Moreover, as demonstrated in Alves et al. (2024), this algorithm is also highly efficient in practice and does not encounter the large constant of fast matrix multiplication. Therefore, Theorem˜1.3 implies that problems with lower bounds based on BMM can be solved more efficiently on structured inputs.
@@
Suppose we are given matrices $\mathbf{M}_{1},\mathbf{M}_{2}\in\{0,1\}^{n\times n}$, where $\mathbf{M}_{i}$ has corrupted VC-dimension $d_{i}$ and $\mathbf{M}_{i}^{\top}$ has corrupted VC dimension $d_{i}^{\prime}$ for $i\in\{1,2\}$. Let $s=\min\{d_{1},d_{2},d_{1}^{\prime},d_{2}^{\prime}\}$. Then, there is an algorithm for Boolean matrix multiplication which runs in $\widetilde{O}(n^{3-1/s})$ time.
@@
This conjecture is used to prove conditional lower bounds for many dynamic graph problems, such as maintaining shortest paths, effective resistances, reachability, bipartite matching, triangle detection, and many others in graphs undergoing vertex insertions and deletions (Henzinger et al., 2015). Even using fast matrix multiplication, no algorithm can beat $O(n^{2})$ time on dense graphs, which is typically equivalent to re-solving them from scratch whenever the graph changes. Since Theorem˜1.3 shows the OMv conjecture does not hold for structured matrices, it leads to improved dynamic algorithms.
Implications to Dynamic Algorithms. We list some novel dynamic algorithms obtained via this technique, whose upper bounds are only achievable due to their structured nature, otherwise violating either the BMM or OMv conjecture. We defer the formal proofs for these results to Appendix˜C.
An important class of linear equations arising in practice have form $\mathbf{L}x=b$, where $\mathbf{L}$ is the Laplacian of an undirected graph $G=(V,E)$ given by $\mathbf{L}\coloneqq\mathbf{D}-\mathbf{A}$ where $\mathbf{D}$ is a diagonal matrix such that $\mathbf{D}_{i,i}=\mathrm{deg}(i)$ for $i\in[n]$ where $|V|=n$, and $\mathbf{A}$ is the adjacency matrix of $G$. There are Laplacian solvers that run in time nearly linear in $|E|$ (Spielman and Teng, 2004). We study the problem of constructing a dynamic Laplacian solver for $\mathbf{L}_{t}x=b_{t}$ where the Laplacian changes over time through vertex updates to the graph, and each $b_{t}$ is an arbitrary new vector.
Previous upper bounds could only handle a $\operatorname{poly}(1/\epsilon)$ accuracy by maintaining spectral sparsifiers (Abraham et al., 2016) or vertex sparsifiers (Durfee et al., 2019). Moreover, exact or $O(\log 1/\epsilon)$-time dependencies are ruled out by the OMv conjecture since no algorithm beats naive recomputation from scratch in $\widetilde{O}(|E|)=\widetilde{O}(n^{2})$ time via nearly-linear time Laplacian solvers (Spielman and Teng, 2004). We give the first result faster than naive recomputation in the high-accuracy $\log(1/\epsilon)$ regime.
@@
With the fast dynamic Laplacian solver, we maintain dynamic effective resistances, which is a critical subroutine in graph clustering (Alev et al., 2017), fault-tolerant computing (Ghosh et al., 2008), network analysis (Tizghadam and Leon-Garcia, 2008; Anand and Umans, 2023; Anand et al., 2024a), and biological systems (Klein et al., 2022), as it measures the connectivity between two vertices. The effective resistance $r_{G}(a,b)$ between vertices $a$ and $b$ in a graph $G$ is a graph-analog of leverage scores that represents the energy needed to route one unit of electric flow from $a$ to $b$. Formally, let $\mathbf{e}_{i}$ denote the $i$-th standard basis vector. Then, for all $a,b\in V$, $r_{G}(a,b)=(\mathbf{e}_{a}-\mathbf{e}_{b})^{\top}\mathbf{L}^{\dagger}(\mathbf{e}_{a}-\mathbf{e}_{b})$, where $\mathbf{L}$ is the Laplacian of $G$ and $\mathbf{L}^{\dagger}$ is the Moore-Penrose pseudoinverse of $\mathbf{L}$.
@@
(Dynamic Effective Resistance). There is a dynamic algorithm that, given a dynamic graph $\smash{G=(V,E)}$ with corrupted VC-dimension bounded by $d$, maintains effective resistances in $G$. The data structure supports queries that receive a pair of vertices $u,v\in V$ and error parameter $\smash{\epsilon>0}$. Then, in $\smash{\widetilde{O}(n^{2-1/d}\log 1/\epsilon)}$ time, the algorithm returns a $\smash{(1\pm\epsilon)}$-approximation of the effective resistance. Moreover, each node update to $G$ in the dynamic data structure takes $\smash{\widetilde{O}(n)}$ time.
Our dynamic matrix-vector multiplication data structure also leads to faster dynamic triangle detection algorithms for graphs with corrupted VC-dimension $d$, which are critical components of fraud detection systems and community-detection algorithms. No subquadratic time algorithm exists for vertex updates, conditional on BMM (Abboud and Vassilevska Williams, 2014) or OMv (Henzinger et al., 2015). We show that these lower bounds can be broken on structured graphs.
@@
(Dynamic Triangle Detection). There is an algorithm that, given a dynamic graph $\smash{G=(V,E)}$ with corrupted VC-dimension $d$, maintains whether $G$ has a triangle or not. Each vertex update takes $\smash{\widetilde{O}(n^{2-1/d})}$ time and returns a Boolean indicator for $G$ containing a triangle.
Our techniques lead to faster dynamic approximate algorithms for single-source distances, where we are tasked with finding the distances from a designated source vertex to every other vertex.
@@
(Dynamic Approximate Single-Source Shortest Paths). There is a dynamic algorithm that maintains $\smash{(1+\epsilon)}$-approximate single-source distances on a dynamic unweighted graph $\smash{G=(V,E)}$. If the corrupted VC-dimension of $G$ is bounded by $d$, each node update to $G$ takes $\smash{\widetilde{O}({kn^{2-1/2d}}/{\epsilon})}$ time, and querying the distances for any source node takes $\smash{\widetilde{O}(n^{2-1/2d}/\epsilon)}$ time.
Given a metric space $\mathcal{M}$ with $n$ points and a positive integer $k\leq n$, the $k$-center problem asks one to select $k$ points, referred to as centers, such that the maximum distance of any point in the metric space to its closest center is minimized. It is known that the $k$-center is NP-hard to $(2-\epsilon)$-approximate for any $\epsilon>0$ (Hsu and Nemhauser, 1979). The best existing dynamic algorithms (Cruciani et al., 2024) only work with edge updates and use fast matrix multiplication as a black-box. We present an algorithm without fast matrix multiplication that supports more powerful vertex updates.
@@
(Dynamic Approximate $k$-center). Given an unweighted undirected graph $\smash{G=(V,E)}$, there is a dynamic algorithm for $\smash{(2+\epsilon)}$-approximate $k$-center with node update time $\smash{\widetilde{O}(kn^{2-1/2d}/\epsilon)}$, where $d$ is a bound on the corrupted VC-dimension of $G$.
@@ 1.3 Further Related Work
Boolean matrix multiplication. A line of work shows that the complexity of matrix-vector multiplication over the Boolean semiring can be mildly subpolynomial. For instance, Williams (2007) showed that BMM can be computed in $O(n^{3}/\log n)$ time. In turn, Larsen and Williams (2017) and Abboud et al. (2024) extended this to $O(n^{3}/2^{\Omega(\sqrt[7]{\log n})})$.
Structured matrices. When the underlying matrix is Vandermonde, Toeplitz, Hankel, or Cauchy, a body of work shows that they admit fast matrix-vector multiplications through convolutional transformations (Motzkin, 1951; Olshevsky and Shokrollahi, 2000; Pan and Tsigaridas, 2014). Orthogonal-vector (OV) matrices are structured matrices where each row and column receives a label $v\in\{0,1\}^{d}$ and $\mathbf{M}_{i,j}=1$ if and only if the corresponding two labels are orthogonal. Similar to our corrupted VC-dimension results, Alman and Vassilevska Williams (2020) show that OV matrices allow fast matrix-vector products if there are few corruptions to the matrix. Matrix products can also be accelerated using geometric data structures Lingas (2002); Floderus et al. (2018). Another approach to handle structured inputs is through the algorithms with predictions regime, where some learning algorithm extracts the structure and can predict information about future vectors, leading to faster matrix-vector multiplication algorithms Henzinger et al. (2024); Brand et al. (2024).
Dynamic structured graphs. Our dynamic algorithms hold for any structured graph, even if we do not know their structure. When the structure is known, there are specialized dynamic algorithms tailored to the specific graph. These include, for instance, interval graphs Crespelle (2019); Chen et al. (2024), geometric graphs defined from Kernels (Alman et al., 2020), dynamic planar graphs (Korhonen et al., 2024; Abboud and Dahlgaard, 2016), and minor free graphs (Dorn et al., 2012).
Matrix vector multiplication in optimization. In convex optimization, there is a long history on developing data structures to accelerate matrix-vector products. Here the matrices are generally projection matrices that require computation of some matrix inverse. While historically, most research was on accelerating the maintenance of the matrix inverse (Karmarkar, 1984; Vaidya, 1989; Lee and Sidford, 2015; Cohen et al., 2021; Lee et al., 2019; Brand, 2020; Jiang et al., 2021, 2020a; Huang et al., 2022; Jiang et al., 2022, 2020b), this research has progressed so far that the current bottleneck for faster linear program solvers are simple matrix-vector products (Brand et al., 2020, 2021).
@@ 2 Preliminaries
Notation. Let $\smash{[n]=\{1,\dots,n\}}$. We use $\smash{O(\cdot)}$ to hide constant factors, and $\smash{\widetilde{O}(\cdot)}$ to hide polylogarithmic factors. Let $\|\cdot\|_{1}$ denote the $\ell_{1}$ (Manhattan distance). For a matrix $\mathbf{M}$, let $\mathbf{M}_{i}$ and $\mathbf{M}_{:,i}$ denote its $i$’th column/row (respectively). For a set $V$ we write $\smash{\mathbf{M}\in\mathbb{R}^{m\times V}}$ for the $\smash{m\times|V|}$ matrix where we can index columns by $\smash{x\in V}$, i.e., $\mathbf{M}_{x}\in\mathbb{R}^{m}$. For any vector $\smash{x\in\mathbb{R}^{n}}$, the Hamming weight of $x$ is the number of non-zero entries of $x$ also denoted by $\operatorname{nnz}(x)$. For Boolean vectors $\smash{x,y\in\{0,1\}^{n}}$, the Hamming distance between $x$ and $y$ is given by $\smash{\|x-y\|_{1}}$. Finally, for any vector $x\in\mathbb{R}^{n}$ and positive definite matrix $\smash{\mathbf{L}\in\mathbb{R}^{n\times n}}$, define the $\mathbf{L}$-induced norm of $x$ by $\smash{\|x\|_{\mathbf{L}}=\sqrt{x^{\top}\mathbf{L}x}}$.
@@
VC-Dimension. A range space (or a set system) is a pair $\mathfrak{R}=(X,R)$ where $X$ is a set and $R$ is a set of subsets of $X$. Let $\Pi_{R}(A)=\{A\cap r:r\in R\}$. The VC-dimension of $\mathfrak{R}$ is $\mathrm{VC}(\mathfrak{R})=\max\{|S|:S\subseteq X\text{ and }|\Pi_{R}(S)|=2^{|S|}\}$. The dual range space of $\mathfrak{R}$ is then given by $\mathfrak{R}^{\top}=(R,\{\{r|x\in r\}|x\in X\})$, and the dual VC-dimension of $\mathfrak{R}$ is the VC-dimension of $\mathfrak{R}^{\top}$. Any matrix $\mathbf{M}\in\{0,1\}^{m\times n}$ corresponds to a set family $\mathcal{F}_{\mathbf{M}}=\{\{j:\mathbf{M}_{i,j}=1\}:i\in[m]\}$, i.e., each row of $\mathbf{M}$ is the indicator vector of some set. Through it, we can encode $\mathbf{M}$ by the range space $\mathfrak{R}_{M}=([n],\mathcal{F}_{\mathbf{M}})$. Then, the VC-dimension of $\mathbf{M}$ is defined as the VC-dimension of $\mathfrak{R}_{\mathbf{M}}$, and the dual VC-dimension is the VC-dimension of $\mathbf{M}^{\top}$. Here, a subset $S$ of the columns of $\mathbf{M}$ is said to be shattered if each of the $2^{|S|}$ many 0/1 strings appears in some row in the restriction of $\mathbf{M}$ to the columns in $S$. Then, the VC-dimension of $\mathbf{M}$ is the maximum size of a shattered subset of the columns of $\mathbf{M}$. Therefore, we see that the VC-dimension $d_{m,n}$ of a matrix $\mathbf{M}\in\{0,1\}^{m\times n}$ satisfies $1\leq d_{m,n}\leq\log m$.